OpenStudy (anonymous):

Factor Completely. (0.01y^2-0.0009)

OpenStudy (anonymous):

how annoying. difference of two squares

OpenStudy (anonymous):

\[a^2-b^2=(a+b)(a-b)\] and here \[a=0.1x,b=0.03\]

OpenStudy (saifoo.khan):

Sateliteeeeeeeeeeeeeeee!

jimthompson5910 (jim_thompson5910):

\[\large 0.01y^2-0.0009\] \[\large \frac{1}{100}y^2-\frac{9}{10000}\] \[\large (\frac{y}{10})^2-(\frac{3}{100})^2\] \[\large (\frac{y}{10}-\frac{3}{100})(\frac{y}{10}+\frac{3}{100})\]

jimthompson5910 (jim_thompson5910):

Alternatively, \[\large 0.01y^2-0.0009\] \[\large (0.1y)^2-(0.03)^2\] \[\large (0.1y-0.03)(0.1y+0.03)\]